OpenStudy (anonymous):

Rename 1/4 and 5/8 using the least common denominator.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

2/8 and 5/8

OpenStudy (anonymous):

yes yes lukeiloklok

OpenStudy (anonymous):

Since LCM of denominators is 8, thus 1/4 * 8= 2 and 5/8* 8= 5/8. The two fractions are therefore 2/8 and 5/8.

OpenStudy (anonymous):

ok lcd im sorry lcd of 1/4 and 5/8 are 2/8 and 5/8

OpenStudy (anonymous):

Rename 3/8 and 6/7 using the least common denominator

OpenStudy (anonymous):

21/56 and 48/56.

OpenStudy (anonymous):

wrong nope

OpenStudy (anonymous):

i think emcrazy is right

OpenStudy (anonymous):

It's correct since LCM of denominators is 56.

OpenStudy (amistre64):

7 (3/8 , 6/7 = N) 8 (3(7)/8 , 6 = 7N) 3(7) , 6(8) = 7(8)N 21 , 48 = 56N 21/56 , 48/56